OpenStudy (saylilbaby):

1. Give the domain, range, intercepts, asymptotes, intervals of increasing and decreasing, intervals of positive and negative, symmetry and extrema (if any) for the graph of this equation. Include a screen shot of the graph.

OpenStudy (saylilbaby):

\[f(x)=\frac{ 3 }{ x^2-4 }+1\]

OpenStudy (amoodarya):

what part of question ,is your problem ?

OpenStudy (saylilbaby):

2. Explain why there are asymptotes for this graph and how to find them algebraically. Use vocabulary given in the text for this module. 3. Algebraically calculate the intercepts, both x and y. Show all work and state extraneous solutions, if any.

OpenStudy (amoodarya):

\[vertical-retrice\rightarrow x^-4=0 \rightarrow x=\pm 2\\ \lim_{x \rightarrow 2}f(x)=\lim_{x \rightarrow 2}\frac{3}{x^2-4}+1=\infty \\so\\x=2 \space is-vertical-as.\\lim_{x \rightarrow -2}f(x)=\lim_{x \rightarrow -2}\frac{3}{x^2-4}+1=\infty \\so\\x=-2 \space is-vertical-as.\] \[\lim_{x \rightarrow \pm \infty}\frac{3}{x^2-4}+1\rightarrow 0+1=1 \rightarrow y=1 \\is\\horizontal-as.\]

OpenStudy (saylilbaby):

can u number em so I know what to put for each number @amoodarya

OpenStudy (amoodarya):

this operation was for finding asymptotes

OpenStudy (amoodarya):

to find intercept find the root(s) of f(x)=0 \[\frac{3}{x^2-4}+1=0\\\frac{3}{x^2-4}=-1\\x^2-4=-3\\x^2=1\\x=\pm1\]